OpenStudy (chupacabraj):

How do you calculate the magnetic force given the velocity and magnetic field vectors?

OpenStudy (chupacabraj):

I need to calculate the magnetic force. q = 2 x 10^-6 C, V= 3i + 10j m/s and B = 0.5i - 0.25k

OpenStudy (chupacabraj):

\[F_{m} = q\left[\begin{matrix}3 & 10 & 0 \\ 0.5 & 0 & -0.25\end{matrix}\right]\]

OpenStudy (chupacabraj):

@agent0smith

OpenStudy (chupacabraj):

Expand cofactor down 2nd column

OpenStudy (chupacabraj):

@zepdrix

OpenStudy (loser66):

determinant is for a square matrix only.

OpenStudy (chupacabraj):

Oh, how am I supposed to calculate the magnetic force then? My teacher says I have to"expand down second column".

OpenStudy (loser66):

I don't know physics

OpenStudy (loser66):

Can you post the original problem?

OpenStudy (chupacabraj):

My professor is just so obsessed with us knowing how to do the cross product by the "expansion of cofactors". He made us do some extra practice problems with finding the determinant of 3 x3 and 4x4 matrices but I now come across this problem, and apparently he wants us to use the same process? PROBLEM: A charged particle has q= + 2micro C. It has velocity v= 3i+10j m/s at a point where B = 0.5i - 0.25k T. What is the magnetic force? (Expand down second column).

OpenStudy (loser66):

oh, cross product is another topic!! :)

OpenStudy (loser66):

the magnetic force \(F_B= q*v*B sin(\theta)\)

OpenStudy (chupacabraj):

Yes, but my teacher wants us to use cross product. He was so pissed the other day when I used a different method. :(

OpenStudy (loser66):

And you need cross product of v and B , right?

OpenStudy (chupacabraj):

yes

OpenStudy (loser66):